Find the equations of the tangent planes to the conicoid 7x^2-3y^2-z^2+21=0, which pass through the line 7x-6y+9=0, z=3
1
Expert's answer
2020-11-17T17:46:52-0500

The equation of the line is "\begin{cases}
7x-6y+9=0
\\
z=3
\end{cases}
\quad
\begin{cases}
x=\frac{6t-9}{7}
\\
y=t
\\
z=3
\end{cases}"

If the plane "ax+by+cz+d=0" passes through this line, then

"a\frac{6t-9}{7}+bt+3c+d=(6a/7+b)t+(3c+d-9a/7)=0"

It means, that "\begin{cases}
6a/7+b=0
\\
3c+d-9a/7=0
\end{cases}
\quad
\begin{cases}
a=7\alpha 
\\
b=-6\alpha 
\\
c=\beta
\\
d=9\alpha-3\beta 
\end{cases}"

So, we have the equation of this plane "7\alpha x-6\alpha y+\beta z+(9\alpha -3\beta)=0"


Let "7\alpha x-6\alpha y+\beta z+(9\alpha -3\beta)=0" be the equation of the tangent plane to the conicoid "7x^2-3y^2-z^2+21=0" at the point "P(x_0,y_0,z_0)" .

The point "P(x_0,y_0,z_0)" is on the conicoid: "7x_0^2-3y_0^2-z_0^2+21=0" .

The equation of the tangent plane is "f_x|_P(x-x_0) +f_y|_P(y-y_0)+f_z|_P(z-z_0)=0" , where "f(x,y,z)=7x^2-3y^2-z^2+21"

We have "14x_0(x-x_0)-6y_0(y-y_0)-2z_0(z-z_0)=0"

"2(7xx_0-3yy_0-zz_0+21)-2(7x_0^2-3y_0^2-z_0^2+21)=0"

Therefore, "7xx_0-3yy_0-zz_0+21=0" is the equation of the tangent plane at "P(x_0,y_0,z_0)" .


Using the fact that "7xx_0-3yy_0-zz_0+21=0" and "7\alpha x-6\alpha y+\beta z+(9\alpha -3\beta)=0" are equations of the same plane, we get

"\begin{cases}
x_0=\alpha 
\\
y_0=2\alpha
\\
z_0=-\beta 
\\
9\alpha -3\beta=21
\end{cases}
\quad
\begin{cases}
x_0=\alpha 
\\
y_0=2\alpha
\\
z_0=7-3\alpha 
\\
\beta=3\alpha -7
\end{cases}"


Substituting "(x_0,y_0,z_0)" into "f(x_0,y_0,z_0)=0" , we get "7\alpha ^2-12\alpha ^2-(3\alpha-7)^2+21=-14\alpha^2+42\alpha -28=0"

There are 2 roots of this equation: "\alpha _1=1" , "\alpha_2 =2" . Then we get "\beta_1=-4" and "\beta_2=-1" .


Therefore, we have two planes: "7x-6y-4z+21=0" and "14x-12y-z+21=0" .


Answer: "7x-6y-4z+21=0" and "14x-12y-z+21=0."